OpenStudy (anonymous):

Can anyone differentiate f(x)= (4sinx)/(2x+cos x). Please show the way to get the answer. Thanks...

OpenStudy (anonymous):

you can use the quotient rule

OpenStudy (anonymous):

Quotient rule with f'(4sin x) = 4 cos x g'(2x + cos x) = 2 - sin x

OpenStudy (anonymous):

Jimmyrep: yes I know but, I still can't get the right answer.

OpenStudy (anonymous):

The answer is (8x cos x - 8sinx + 4) / (2x + cos x)^2

OpenStudy (anonymous):

Are you okay with the denominator?

OpenStudy (anonymous):

Did you see my derivatives: f'g - g'f (4cosx)(2x + cosx) - (2 - sin x)(4 sin x)

OpenStudy (anonymous):

(2x + cos x)* 4 cos x - 4 sinx * (2 - sin x) / (2x + cos x)^2 = 8x cos x + 4 cos^2x - 8 sin x + 4 sin^2 x / (2x + cos x)^2 =[ 8 (x cos x - sin x) + 4 ] / (2x + cos x)^2

OpenStudy (anonymous):

yes i got the step before the final answer. but I still do not understand how get their answer. help please.

OpenStudy (anonymous):

sin^2x + cos^x = 1 so 4(cosx^x + sin^x) = 4

OpenStudy (anonymous):

oh thanks.. i remember that use identities of trigonometry.. thank you very much all.

OpenStudy (anonymous):

yw